OpenStudy (anonymous):

Mr. Sato brought 5 dozen hot dog buns priced at 6 buns for $1.29 and 6 packages of hotdogs priced at 2 packages for $5.50. What is the total amount he spent , not including tax, on the buns and hot dogs?

OpenStudy (lilg132):

oh misread the question

OpenStudy (anonymous):

i know right! Its so confusing!

OpenStudy (lilg132):

arghhhh made a mistake

OpenStudy (anonymous):

lol

OpenStudy (lilg132):

5 * 12 = 60 60 / 6 = 10 10 * 1.29 = 12.9 6 / 2 = 3 3 * 5.50 = 16.5 12.9 + 16.5 = $29.4 DONT FORGET TO CLICK GOOD ANSWER :D